OpenStudy (anonymous):

jim is able to sell a hand carved statue for $670 which was a 35% profit over his cost. How much did his statue originally cost?

OpenStudy (anonymous):

\[670=x+0.35x\] solving the equation for x gives: \[x=496.3\]

OpenStudy (anonymous):

so it costs $496.3

OpenStudy (anonymous):

how did u solve for x what did u do show me

OpenStudy (anonymous):

well x+0.35x=670 => 1.35x=670 x=670/1.35=496.3 make sense?

OpenStudy (anonymous):

yes i have to do all of this without a calculator

OpenStudy (anonymous):

you should

OpenStudy (anonymous):

do you know how I got the equation in the first place?

OpenStudy (anonymous):

no see i have a learning disability and sometimes things are a little complicating for me to understand.

OpenStudy (anonymous):

nah you don't.. it's as you said that 670 is the original cost plus 35% profit.. so considering x the original cost we will have: 670=x+35%(x) which is the same as 670=x+0.35x I hope it's clear

OpenStudy (anonymous):

ok i got it now thank you so much :D

OpenStudy (anonymous):

you're welcome :)